# Teleportation=Translation: Continuous recovery of black hole information

### v3

@@ Abstract
[…]
@@ 1 Introduction
The black hole information paradox exposes a fundamental tension between the apparent thermality of Hawking radiation, derived from semiclassical gravity, and the unitary evolution required by quantum theory Hawking1976 . Although Hawking’s original calculation suggests that a semiclassical black hole evaporates via thermal radiation in a manner violating unitarity, recent developments indicate that information recovery is possible even within the semiclassical framework, provided the paradox is analyzed with refined conceptual and technical tools Almheiri2020 ; Raju2022 ; Akil:2025bhip .
@@
In particular, the AdS/CFT correspondence has provided a semiclassical framework utilizing replica wormholes and quantum extremal surfaces to derive the Page curve Almheiri2019 ; Almheiri2020Page ; Penington2022 . While these results strongly support the preservation of unitarity, they raise critical questions regarding the interpretation of the gravitational path integral as an ensemble average SaadShenkerStanford and its compatibility with spacetime factorization HarlowJafferis2020 ; Marolf2020 . To address these foundational issues, quantum information and operator-algebraic approaches have been increasingly employed to clarify the underlying assumptions behind these entropy computations Raju2022 ; Kibe2022 .
@@
However, calculating the entropy curve does not, by itself, explain the dynamical mechanism of information retrieval. To focus on the retrieval process, Hayden and Preskill reformulated the problem as a quantum decoding task. They demonstrated that by modeling the black hole as a fast scrambler, any new information falling in can be retrieved rapidly, provided the observer has access to the early radiation HaydenPreskill . This recovery process is operationally analogous to quantum teleportation. Building on this operational viewpoint, van den Heijden and Verlinde (vdH-V) reformulated the problem entirely within the language of operator algebras vdHV . Mathematical foundations for such teleportation protocols have been developed in Conlon2022 . In their framework, the canonical shift $\Gamma$ plays the role of a teleportation operator, leading to the identification Teleportation=Translation.
@@
Nonetheless, as acknowledged in vdHV , a direct realization of this protocol in the continuum limit faces fundamental obstructions. Local operator algebras in quantum field theory are Type III von Neumann factors, distinguished by the absence of a tracial state Connes80 ; Witten . This algebraic feature reflects the physical reality that vacuum entanglement is effectively unbounded. Consequently, the maximally entangled resources prerequisite for standard teleportation are mathematically ill-defined in this setting.
@@
In this work, we construct a smooth unitary interpolation to connect the discrete canonical shift with continuous spacetime evolution. A central challenge in this construction is ensuring that this interpolation path is canonical rather than arbitrary. To resolve this in the Type III context, we employ the Haagerup–Kosaki framework Connes80 ; Kosaki84 and lift the inclusion to a Type II crossed product via Takesaki duality Connes80 ; TakesakiII . Crucially, the conditional expectation in this setting is uniquely determined by the invariance of the reference weights defining the $L^{p}$-interpolation Kosaki84 ; Takesaki72 . This rigidity guarantees that our unitary path is intrinsic to the modular structure of the system. Furthermore, by exploiting the analytic properties of modular flow, we ensure the path is sufficiently smooth to define a continuous generator $\widetilde{G}$ as the unique self-adjoint operator associated with this flow.
@@
Building on this construction, we derive our main result by proving that this generator bears a direct physical identification with the modular momentum. Specifically, we establish an exact operator identity asserting that the generator of the teleportation flow equals twice the difference of the modular Hamiltonians ($\widetilde{G}=2P$). The factor of 2 is dictated by Borchers’ commutation relations Borchers92 , ensuring consistency with the scaling of modular flows. Our proof establishes this equality on a dense core of analytic modular vectors and demonstrates the equality of the closed operators via Nelson’s analytic vector theorem Nelson59 . Furthermore, we quantify the stability of this identification under small deformations using non-commutative $L^{p}$-theory Kosaki84 ; Kato . On a physical level, this result provides a precise operator-algebraic realization of the Teleportation=Translation proposal, demonstrating that information recovery manifests as a continuous geometric translation. Finally, to facilitate the verification of this operator identity in complex physical models where direct analysis is challenging, we propose a concrete observable test based on correlation functions.
@@
The remainder of this paper is organized as follows. In Sec. 2, we summarize the necessary elements of Tomita–Takesaki theory and introduce the Jones projection together with the canonical shift $\Gamma$. In Sec. 3, we construct the continuous unitary interpolation by lifting the Type III inclusion to the crossed product via the Haagerup–Kosaki framework. Sec. 4 is devoted to the proof of our main result, identifying the generator of this flow with the modular momentum ($\widetilde{G}=2P$). We conclude with a discussion of the results in Sec. 5.
@@ 2.1 Preliminaries: Tomita–Takesaki Theory
We work on a fixed Gelfand-Naimark-Segal (GNS) Hilbert space $\mathcal{H}$ built from a faithful normal state $\omega$ (with cyclic and separating vector $|\Omega\rangle$) for the von Neumann algebra $\mathcal{M}$ TakesakiI . Let $\mathcal{N}\subset\mathcal{M}$ be an inclusion of von Neumann algebras. For an algebra $\mathcal{X}$ (where $\mathcal{X}=\mathcal{M}$ or $\mathcal{N}$), the core objects of Tomita–Takesaki theory TakesakiII are:
@@
  • The modular Hamiltonian $K_{\mathcal{X}}$, defined by $\Delta_{\mathcal{X}}=e^{-2\pi K_{\mathcal{X}}}$.
@@ 2.2 The Canonical Shift Protocol
Step 1: Conditional Expectation. Let $E:\mathcal{M}\to\mathcal{N}$ be a normal, faithful conditional expectation that is $\omega$-preserving: $\omega(E(a))=\omega(a)$ for all $a\in\mathcal{M}$. This map effectively discards information in $\mathcal{M}$ that is not in $\mathcal{N}$.
@@
$E$ being a $\omega$-preserving conditional expectation ensures that $e_{\mathcal{N}}$ is a well-defined self-adjoint projection ($e_{\mathcal{N}}^{2}=e_{\mathcal{N}}$, $e_{\mathcal{N}}^{\dagger}=e_{\mathcal{N}}$) and satisfies $e_{\mathcal{N}}ae_{\mathcal{N}}=E(a)e_{\mathcal{N}}$ for all $a\in\mathcal{M}$. The discrete (0 or 1) nature of this projection is the source of the protocol’s discreteness.
@@
Information is not lost, but relocated from one relative commutant to another. Crucially, while explicitly defined algebraically, this shift operator $U_{\Gamma}$ corresponds to a geometric translation in the emergent spacetime—a connection that will be precisely established in the continuum limit in the subsequent sections.
@@ 3 From a Discrete Protocol to a Continuous Path
Our core strategy is to soften the discrete Jones projection $e_{\mathcal{N}}$ by constructing a continuous family of maps $E_{s}$ and corresponding operators $e_{\mathcal{N}}(s)$ for $s\in[0,1]$, such that $E_{0}=\mathrm{id}$ and $E_{1}=E$. We first briefly discuss the challenge of (dynamical) idempotency in Sec. 3.1. Then, we present the pedagogical completely positive interpolation (Path A). Although this path ultimately fails the idempotency requirement in the infinite-dimensional (Type III) setting, discussing it is crucial for understanding the necessity of the more advanced construction (Sec. 3.2). We then lift the problem to the semifinite crossed-product (Haagerup–Kosaki envelope) and construct the canonical, physical interpolating path via $L^{p}$-space interpolation theory (Path B), satisfying the rigorous boundary conditions required for our proof (Sec. 3.3). Finally, using the Tomita–Takesaki machinery in the semifinite setting, we build the unitary path $\widetilde{U}(s)$ and analytically establish the generator $\widetilde{G}$ as a tangent vector, proving its self-adjointness via analytic vectors (Sec. 3.4).
@@ 3.1 The Physical Requirement: Dynamical Idempotency
To define a continuous unitary path $U(s)=J_{\mathcal{M}}J_{\mathcal{N}(s)}$, we must first construct a continuous path of von Neumann subalgebras $\mathcal{N}(s)$ such that $\mathcal{N}(0)=\mathcal{M}$ and $\mathcal{N}(1)=\mathcal{N}$. Such a path of algebras would, in turn, define a path of conditional expectations $E_{s}:\mathcal{M}\to\mathcal{N}(s)$. The crucial link between the map $E_{s}$ and the subalgebra $\mathcal{N}(s)$ is idempotency. Furthermore, for the parameter $s$ to be interpreted as a physical coarse-graining or information resolution flow, one should expect a condition even stronger than simple “static” idempotency ($E_{s}^{2}=E_{s}$). A true physical flow should satisfy a “dynamical” idempotency or semigroup property, such as $E_{s^{\prime}}\circ E_{s}=E_{s^{\prime}}$ for $s^{\prime}\geq s$. This condition, common in processes like the renormalization group, ensures the path is consistent: coarse-graining to level $s$ and then further to $s^{\prime}$ is identical to coarse-graining directly to $s^{\prime}$. While this dynamical property implies the static one, the static property ($E_{s}^{2}=E_{s}$) remains the minimum mathematical requirement for $E_{s}$ to be a conditional expectation at all, as established by the following theorem (Tomiyama theorem Tomiyama57 ).
@@
Dynamic idempotency for a family of maps $\{E_{s}\}_{s\in I}$ means:
@@
$E_{s^{\prime}}\circ E_{s}=E_{s^{\prime}}\quad\text{for }s^{\prime}\geq s\qquad\text{(semigroup/absorption property)}.$ (4)
While static idempotency is sufficient to define a projection at a single point, the dynamical condition is essential for interpreting the path as a coherent teleportation protocol. Physically, this ensures that the sequence of coarse-graining operations is consistent: projecting to an intermediate scale $s$ and then to a coarser scale $s^{\prime}$ must be identical to projecting directly to $s^{\prime}$. In the context of information recovery, this implies a strict nesting of information. Without this property, the intermediate stages of the protocol would lack a hierarchical structure, meaning that recovered information at step $s$ could be inexplicably lost or scrambled in subsequent steps. Thus, dynamical idempotency is the rigorous condition that prevents information leakage and ensures the monotonicity of the recovery flow from the black hole interior ($\mathcal{N}$) to the exterior ($\mathcal{M}$).
Theorem 3.2.
(Idempotent CP maps are conditional expectations) Let $\mathcal{M}$ be a von Neumann algebra and $\omega$ a faithful normal state on $\mathcal{M}$. Let $E:\mathcal{M}\to\mathcal{M}$ be a normal, unital, completely positive (CP) map satisfying $E^{2}=E$ and $\omega\circ E=\omega$. Then $\mathcal{N}:=E(\mathcal{M})$ is a von Neumann subalgebra of $\mathcal{M}$, and $E$ is the faithful normal conditional expectation from $\mathcal{M}$ onto $\mathcal{N}$.
This theorem sets our target: we must find a continuous path of idempotent CP maps.
3.2 Path A: Pedagogical (CP-Map) Interpolation
A first, intuitive attempt to connect the identity map ($\mathrm{id}$) to the target conditional expectation ($E$) is to construct a simple linear interpolation. The primary requirement is that every intermediate map must be a valid quantum channel, i.e., a unital CP map.
The most direct construction is a convex combination of the maps themselves:
@@ 3.2 Path A: Pedagogical (CP-Map) Interpolation
Since both $\mathrm{id}$ and $E$ are unital CP maps, their convex combination $E_{s}$ is automatically a unital CP map for all $s$, regardless of the underlying algebra type.
  • Finite Dimensions: This corresponds to linearly interpolating the Choi matrices Choi75 , preserving positivity throughout the path.
  • Infinite Dimensions: For general von Neumann algebras (including Type III), Stinespring’s dilation theorem Stinespring55 guarantees that any such CP map $E_{s}$ can be physically realized via an isometry $V_{s}$ on a larger Hilbert space (i.e., $E_{s}(x)=V_{s}^{\dagger}\pi(x)V_{s}$).
These constructions provide a continuous path of maps $E_{s}$ that satisfy the fundamental mathematical properties of well-definedness, boundedness, and self-adjointness. We establish these properties first for the map $E_{s}$, and then for the corresponding GNS operator $e_{\mathcal{N}}(s)$.
@@
    $\omega$-preserving, i.e., $\omega(E_{s}(a))=\omega(a)$ for all $a\in\mathcal{M}$.
@@
    Self-adjoint with respect to the GNS inner product associated with $\omega$.
@@
(1) State-preserving: Both $\mathrm{id}$ and $E$ are $\omega$-preserving (by assumption for $E$). Since $E_{s}$ is constructed as a convex combination (at the level of maps or their Choi/Stinespring representations), the property follows by linearity:
$\omega(E_{s}(a))=(1-s)\omega(\mathrm{id}(a))+s\,\omega(E(a))=\omega(a).$ (6)
(2) GNS Self-adjointness: A map $T$ is self-adjoint with respect to the GNS inner product if $\omega(y^{*}T(x))=\omega((T(y))^{*}x)$ for all $x,y\in\mathcal{M}$. This property holds for both $\mathrm{id}$ and the $\omega$-preserving conditional expectation $E$. By linearity, it extends to $E_{s}$:
@@
This lemma allows us to prove the properties of the GNS operator.
@@
(Existence, Boundedness, and Self-Adjointness of $e_{\mathcal{N}}(s)$) For each $s\in[0,1]$, the operator $e_{\mathcal{N}}(s)$ defined on the dense domain $D=\{|a\rangle\mid a\in\mathcal{M}\}\subset\mathcal{H}$ by
@@
extends uniquely to a bounded, self-adjoint linear operator on $\mathcal{H}$ with operator norm $\|e_{\mathcal{N}}(s)\|\leq 1$. Furthermore, it satisfies the boundary conditions $e_{\mathcal{N}}(0)=\mathbf{1}$ and $e_{\mathcal{N}}(1)=e_{\mathcal{N}}$.
@@
(1) Boundedness and Well-definedness: Since $E_{s}=(1-s)\mathrm{id}+sE$ is a convex combination of unital CP maps (the identity and the conditional expectation), $E_{s}$ itself is a unital CP map. We utilize the Kadison-Schwarz inequality for unital positive maps, which states $E_{s}(a)^{*}E_{s}(a)\leq E_{s}(a^{*}a)$ for all $a\in\mathcal{M}$. Evaluating this in the state $\omega$ and using the property that $\omega\circ E_{s}=\omega$ (since both $\mathrm{id}$ and $E$ preserve $\omega$), we find:
$\||E_{s}(a)\rangle\|^{2}$ $=\omega(E_{s}(a)^{*}E_{s}(a))$
$\leq\omega(E_{s}(a^{*}a))$
$=\omega(a^{*}a)=\||a\rangle\|^{2}.$ (9)
This inequality $\|e_{\mathcal{N}}(s)|a\rangle\|\leq\||a\rangle\|$ has two implications:
  • Well-definedness: If $|a\rangle=0$ (i.e., $\omega(a^{*}a)=0$), then the inequality implies $\||E_{s}(a)\rangle\|=0$, so $|E_{s}(a)\rangle=0$. Thus, the mapping $|a\rangle\mapsto|E_{s}(a)\rangle$ respects the equivalence classes of the GNS construction and defines a valid operator on $D$.
  • Boundedness: The operator is bounded on the dense domain $D$ with $\|e_{\mathcal{N}}(s)\|\leq 1$. By the bounded linear transformation (B.L.T.) theorem, it extends uniquely to a bounded linear operator defined on the entire Hilbert space $\mathcal{H}$.
(2) Self-adjointness: Consider the sesquilinear form associated with $e_{\mathcal{N}}(s)$ on the domain $D$:
$\langle|a\rangle|e_{\mathcal{N}}(s)|b\rangle\rangle=\langle|a\rangle||E_{s}(b)\rangle\rangle=\omega(a^{*}E_{s}(b)).$ (10)
Since $E_{s}$ is a real linear combination of self-adjoint maps (the identity map and the conditional expectation $E$, both of which satisfy $\omega(x^{*}T(y))=\omega(T(x)^{*}y)$), $E_{s}$ is also self-adjoint with respect to the inner product induced by $\omega$. Explicitly:
$\langle|a\rangle|e_{\mathcal{N}}(s)|b\rangle\rangle$ $=(1-s)\omega(a^{*}b)+s\,\omega(a^{*}E(b))$
$=(1-s)\omega(a^{*}b)+s\,\omega(E(a)^{*}b)\quad(\text{property of }E)$
$=\omega(((1-s)a+sE(a))^{*}b)$
$=\omega(E_{s}(a)^{*}b)=\langle e_{\mathcal{N}}(s)|a\rangle||b\rangle\rangle.$ (11)
Thus, $e_{\mathcal{N}}(s)$ is a symmetric bounded operator defined on the entire Hilbert space $\mathcal{H}$. For bounded operators defined everywhere, symmetry is equivalent to self-adjointness.
(3) Boundary Conditions: By definition, $E_{0}=\mathrm{id}$, so $e_{\mathcal{N}}(0)|a\rangle=|a\rangle$, implying $e_{\mathcal{N}}(0)=\mathbf{1}$. Similarly, $E_{1}=E$, so $e_{\mathcal{N}}(1)|a\rangle=|E(a)\rangle$, which corresponds to the Jones projection $e_{\mathcal{N}}$. ∎
These paths, therefore, define continuous families of self-adjoint operators that are mathematically well-behaved. However, they suffer from a fundamental defect: for intermediate values $0<s<1$, the interpolated maps fail to be idempotent,
$E_{s}^{2}\neq E_{s}\quad\text{for }s\in(0,1).$ (12)
As a result, the image $E_{s}(\mathcal{M})$ does not form a von Neumann subalgebra, and Theorem 3.2 cannot be applied. In particular, the modular structures associated with $E_{s}(\mathcal{M})$, such as the conjugation $J_{E_{s}(\mathcal{M})}$, are ill-defined for these intermediate steps, rendering the path physically incomplete.

### v4

@@ Abstract
The Teleportation=Translation conjecture posits that the recovery of information from a black hole is dual to a geometric translation in the emergent spacetime. In this paper, we establish this equivalence for general local quantum field theories by constructing a continuous unitary interpolation that bridges discrete algebraic teleportation protocols and continuous modular flow. We resolve the failure of dynamic idempotency, fundamentally inherent in Type III von Neumann algebras, by employing the Haagerup-Kosaki crossed-product construction. This lift to the semifinite Type II envelope yields a canonical, dynamically consistent path. Crucially, we prove that its unique infinitesimal generator $\tilde{G}$ is exactly twice the geometric modular momentum ($\tilde{G}=2P$). We establish this identity as a closed operator equivalence using Nelson’s analytic vector theorem and quantify its structural robustness via non-commutative $L^{p}$ theory. Ultimately, our results demonstrate that unitary information recovery fundamentally manifests as a continuous geometric translation. This provides a rigorous operator-algebraic mechanism for resolving the black hole information paradox, offering a kinematic framework naturally extendable to include gravitational back-reaction.
@@ 1 Introduction
The black hole information paradox exposes a fundamental tension between the apparent thermality of Hawking radiation, derived from semiclassical gravity, and the unitary evolution required by quantum theory Hawking1976 . Although Hawking’s original calculation suggested that a semiclassical black hole evaporates via thermal radiation in a manner violating unitarity, recent developments indicate that information recovery is possible even within the semiclassical framework, provided the problem is analyzed with refined conceptual and technical tools Almheiri2020 ; Raju2022 ; Akil:2025bhip .
@@
In particular, the AdS/CFT correspondence has provided a semiclassical framework utilizing replica wormholes and quantum extremal surfaces to derive the Page curve Almheiri2019 ; Almheiri2020Page ; Penington2022 . While these results strongly support the preservation of unitarity, they raise critical questions regarding the interpretation of the gravitational path integral as an ensemble average SaadShenkerStanford and its compatibility with spacetime factorization HarlowJafferis2020 ; Marolf2020 . To address these foundational issues, quantum information and operator-algebraic approaches have been increasingly employed to clarify the underlying assumptions behind these entropy computations Raju2022 ; Almheiri_2015 ; Penington19 ; Kibe2022 .
@@
However, calculating the entropy curve does not, by itself, explain the dynamical mechanism of information retrieval. To address the retrieval process, Hayden and Preskill reformulated the problem as a quantum decoding task, demonstrating that a black hole acting as a fast scrambler allows for rapid information retrieval, provided the observer has access to the early radiation HaydenPreskill . This recovery process is operationally analogous to quantum teleportation. Building on this operational viewpoint, van den Heijden and Verlinde (vdH-V) reformulated the problem entirely within the language of operator algebras vdHV , with mathematical foundations for such protocols further developed in Conlon2022 . In their framework, passing to the continuum limit transforms the canonical shift into a continuous teleportation operator. Furthermore, they demonstrated that in the special symmetric case of a half-sided modular inclusion (HSMI), this naturally leads to the identification Teleportation = Translation.
@@
Nonetheless, while the continuous limit naturally emerges under the special symmetric conditions of a HSMI, extending this protocol to the continuum limit for general local quantum field theories faces fundamental obstructions, as acknowledged in vdHV . Local operator algebras in quantum field theory are Type III von Neumann factors, distinguished by the absence of a tracial state Connes80 ; Witten . This algebraic feature reflects the physical reality that vacuum entanglement is effectively unbounded. Consequently, the maximally entangled resources prerequisite for standard teleportation are mathematically ill-defined in this general setting.
@@
In this work, we generalize this paradigm to arbitrary local quantum field theories. We construct a mathematically rigorous and physically consistent smooth unitary interpolation that connects the discrete canonical shift with continuous spacetime evolution. A central challenge is ensuring that this interpolation path is canonical rather than arbitrary. To resolve this in the general Type III context, we employ the Haagerup–Kosaki framework Connes80 ; Kosaki84 and lift the inclusion to a Type II crossed product via Takesaki duality Connes80 ; TakesakiII . Crucially, the conditional expectation in this setting is uniquely determined by the invariance of the reference weights defining the $L^{p}$-interpolation Kosaki84 ; Takesaki72 . This rigidity guarantees that our unitary path is intrinsic to the modular structure of the system. Furthermore, by exploiting the analytic properties of modular flow, we ensure the path is sufficiently smooth to rigorously define the infinitesimal generator $\widetilde{G}$ at the interpolation boundary $s=0$ as the unique self-adjoint operator driving this flow.
@@
Building on this construction, we derive our main result: the generator of the teleportation flow bears a direct physical identification with the modular momentum. Specifically, we establish an exact operator identity asserting that the infinitesimal generator equals twice the difference of the modular Hamiltonians, $\widetilde{G}=2P$. The factor of 2 is dictated by Borchers’ commutation relations Borchers92 , ensuring consistency with the scaling of modular flows. Our proof establishes this equality on a dense core of analytic modular vectors and extends it to the closed operators via Nelson’s analytic vector theorem Nelson59 . Moreover, we quantify the stability of this identification under small deformations using non-commutative $L^{p}$-theory Kosaki84 ; Kato .
On a physical level, this result directly addresses the ultimate motivation of this work: the black hole information paradox. It demonstrates that infalling information is not lost, but rather unitarily recovered through a continuous path. From an algebraic viewpoint, this recovery process constitutes quantum teleportation, which our exact operator identity proves is dynamically equivalent to a geometric translation in spacetime. By rigorously establishing this Teleportation = Translation identity for general Type III inclusions—without relying on the symmetries of a HSMI—our work confirms that information recovery fundamentally manifests as a continuous geometric translation. Finally, to facilitate verification in complex physical models, we propose a straightforward test based on two-point correlation functions.
@@
The remainder of this paper is organized as follows. In Sec. 2, we summarize the necessary elements of Tomita–Takesaki theory and introduce the Jones projection together with the canonical shift. In Sec. 3, we construct the continuous unitary interpolation by lifting the Type III inclusion to the crossed product via the Haagerup–Kosaki framework. Sec. 4 is devoted to the proof of our main result, identifying the generator of this flow with the modular momentum ($\widetilde{G}=2P$). We conclude with a discussion of the results in Sec. 5.
@@ 2.1 Preliminaries: Tomita–Takesaki Theory
We work on a fixed Gelfand-Naimark-Segal (GNS) Hilbert space $\mathcal{H}$ TakesakiI built from a faithful normal state $\omega$ (with cyclic and separating vector $|\Omega\rangle$) for the von Neumann algebra $\mathcal{X}$. The core objects of Tomita–Takesaki theory TakesakiII are:
@@
  • The modular Hamiltonian $K_{\mathcal{X}}$, defined by $\Delta_{\mathcal{X}}=e^{-K_{\mathcal{X}}}$.
@@ 2.2 The Canonical Shift Protocol
Step 1: Conditional Expectation. Let $\mathcal{N}\subset\mathcal{M}$ be an inclusion of von Neumann algebras and $E:\mathcal{M}\to\mathcal{N}$ be a normal, faithful conditional expectation that is $\omega$-preserving: $\omega(E(a))=\omega(a)$ for all $a\in\mathcal{M}$. This map effectively discards information in $\mathcal{M}$ that is not in $\mathcal{N}$.
@@
$E$ being a $\omega$-preserving conditional expectation ensures that $e_{\mathcal{N}}$ is a well-defined self-adjoint projection ($e_{\mathcal{N}}^{2}=e_{\mathcal{N}}$, $e_{\mathcal{N}}^{\dagger}=e_{\mathcal{N}}$) and satisfies $e_{\mathcal{N}}ae_{\mathcal{N}}=E(a)e_{\mathcal{N}}$ for all $a\in\mathcal{M}$.
@@
Information is not lost, but relocated from one relative commutant to another. This suggests that the algebraic shift operator $U_{\Gamma}$ acts as a geometric translation that maps the interior sector $\mathcal{N}^{\prime}\cap\mathcal{M}$ to the radiation sector $\mathcal{M}^{\prime}\cap\mathcal{M}_{1}$. This identification between algebraic relocation and spacetime translation will be established precisely in the continuum limit in the subsequent sections.
@@ 3 From a Discrete Protocol to a Continuous Path
Our core strategy is to construct a continuous teleportation flow. To this end, we first discuss the physical requirement of dynamic idempotency in Sec. 3.1. In Sec. 3.2, we present a pedagogical interpolation (Path A), which, despite satisfying several key physical properties, ultimately fails to maintain idempotency. To resolve this, we lift the problem to the semifinite crossed-product (Haagerup–Kosaki envelope) and construct a canonical physical interpolating path via non-commutative $L^{p}$-space interpolation theory (Path B), which rigorously satisfies the idempotency requirement (Sec. 3.3). Finally, by employing Tomita–Takesaki machinery in this semifinite setting, we build a unitary path $\tilde{U}(s)$ and analytically establish its generator $\tilde{G}$ as the tangent vector at $s=0$, proving its self-adjointness through the theory of analytic vectors (Sec. 3.4).
@@ 3.1 The Physical Requirement: Dynamical Idempotency
To define a continuous unitary path $U(s)=J_{\mathcal{M}}J_{\mathcal{N}(s)}$ for $s\in[0,1]$, we must first construct a continuous family of von Neumann subalgebras $\mathcal{N}(s)$ interpolating between the full algebra $\mathcal{N}(0)=\mathcal{M}$ and the target subalgebra $\mathcal{N}(1)=\mathcal{N}$. This path of algebras is operationally characterized by a corresponding family of conditional expectations $E_{s}:\mathcal{M}\to\mathcal{N}(s)$, where the connection between the map $E_{s}$ and the subalgebra $\mathcal{N}(s)$ is fundamentally established by the property of idempotency.
However, for the parameter $s$ to be physically interpreted as a coarse-graining scale or an information resolution flow, a condition stronger than simple static idempotency ($E_{s}^{2}=E_{s}$) is required. A consistent physical flow must satisfy dynamic idempotency (or the semigroup property): $E_{s^{\prime}}\circ E_{s}=E_{s^{\prime}}$ for $s^{\prime}\geq s$. This condition ensures that the sequence of coarse-graining operations is consistent—strictly analogous to the renormalization group flow—and implies a strict nesting of information resources. Mathematically, this prevents information leakage and guarantees that the image of the map at every step retains the structure of a von Neumann algebra. Consequently, it ensures the monotonicity of the information recovery flow from the black hole interior to the exterior.
@@
Dynamic idempotency for a family of maps $\{E_{s}\}_{s\in[0,1]}$ is defined by the following conditions:
@@
$E_{s^{\prime}}\circ E_{s}=E_{s^{\prime}}\quad\text{for }s^{\prime}\geq s\qquad\text{(semigroup property)}.$ (4)
By Tomiyama’s theorem Tomiyama57 , any normal, unital, completely positive (CP) map satisfying static idempotency $E^{2}_{s}=E_{s}$ (along with state invariance) constitutes a faithful normal conditional expectation onto the von Neumann subalgebra $\mathcal{N}(s):=E_{s}(\mathcal{M})$. Our objective is thus reduced to constructing a continuous path of such idempotent CP maps that further satisfies the dynamic consistency requirement.
3.2 Path A: Pedagogical Interpolation
The most intuitive approach to connecting the identity map ($\mathrm{id}$) to the target conditional expectation ($E$) is to construct a linear interpolation. The primary constraint for any physical process is that every intermediate map must constitute a valid quantum channel—that is, a unital CP map.
The most direct construction satisfying this requirement is the convex combination of the endpoint maps:
@@ 3.2 Path A: Pedagogical Interpolation
Since both $\mathrm{id}$ and $E$ are unital CP maps, their convex combination $E_{s}$ inherently preserves these properties. Thus, $E_{s}$ remains a well-defined quantum channel for all $s$, even in the Type III setting.
  • Finite Dimensions: In matrix algebras, this construction is equivalent to linearly interpolating the corresponding Choi matrices Choi75 . Since the Choi matrices of the endpoint maps $\mathrm{id}$ and $E$ are positive semidefinite, their convex combination preserves this property, ensuring complete positivity along the entire path.
  • Infinite Dimensions: For general von Neumann algebras (including Type III factors), Stinespring’s dilation theorem Stinespring55 guarantees that every intermediate map $E_{s}$ admits a physical realization. Specifically, there exists a dilation Hilbert space $\mathcal{K}_{s}$, a representation $\pi_{s}:\mathcal{M}\to\mathcal{B}(\mathcal{K}_{s})$, and an isometry $V_{s}:\mathcal{H}\to\mathcal{K}_{s}$ such that the map acts as a compression:
    $E_{s}(x)=V_{s}^{\dagger}\pi_{s}(x)V_{s},\quad\forall x\in\mathcal{M}.$
These constructions provide a continuous path of maps $E_{s}$ that satisfy the fundamental mathematical properties of is mathematically well-definedness, boundedness, and self-adjointness. We first establish these properties for the map $E_{s}$ itself, and subsequently for the corresponding operator $e_{\mathcal{N}(s)}$ on the GNS Hilbert space.
@@
    $\omega$-preserving: $\omega(E_{s}(a))=\omega(a)$ for all $a\in\mathcal{M}$.
@@
    Self-adjoint with respect to the GNS inner product associated with $\omega$: $\omega(y^{*}E_{s}(x))=\omega((E_{s}(y))^{*}x)$ for all $x,y\in\mathcal{M}$.
@@
(1) State-preserving: By assumption, $\omega\circ E=\omega$. Since $E_{s}$ is a linear combination,
$\omega(E_{s}(a))=(1-s)\omega(a)+s\,\omega(E(a))=\omega(a).$ (6)
(2) GNS Self-adjointness: The condition for self-adjointness with respect to the GNS inner product holds for both $\mathrm{id}$ and the $\omega$-preserving conditional expectation $E$. By linearity,
@@
Using this lemma, we can rigorously define the interpolated operator on the Hilbert space.
@@
(Existence and Properties of $e_{\mathcal{N}}(s)$) For each $s\in[0,1]$, the operator $e_{\mathcal{N}}(s)$ defined on the dense domain $D=\{|a\rangle\mid a\in\mathcal{M}\}\subset\mathcal{H}$ by
@@
extends uniquely to a bounded, self-adjoint linear operator on $\mathcal{H}$ with norm $\|e_{\mathcal{N}}(s)\|\leq 1$. Furthermore, it satisfies the boundary conditions $e_{\mathcal{N}(0)}=\mathbf{1}$ and $e_{\mathcal{N}(1)}=e_{\mathcal{N}}$.
@@
(1) Boundedness: Using the Kadison-Schwarz inequality for the unital CP map $E_{s}$, we have $E_{s}(a)^{*}E_{s}(a)\leq E_{s}(a^{*}a)$. Evaluating in the state $\omega$ (which is preserved by $E_{s}$):
$\||E_{s}(a)\rangle\|^{2}=\omega(E_{s}(a)^{*}E_{s}(a))\leq\omega(E_{s}(a^{*}a))=\omega(a^{*}a)=\||a\rangle\|^{2}.$ (9)
This implies well-definedness (respecting GNS equivalence classes) and boundedness ($\|e_{\mathcal{N}(s)}\|\leq 1$), allowing a unique extension to $\mathcal{H}$.
(2) Self-adjointness: The sesquilinear form is symmetric due to the self-adjointness of $E_{s}$ (Lemma 3.2):
$\langle|a\rangle|e_{\mathcal{N}(s)}|b\rangle\rangle=\omega(a^{*}E_{s}(b))=\omega(E_{s}(a)^{*}b)=\langle e_{\mathcal{N}(s)}|a\rangle||b\rangle\rangle.$ (10)
For a bounded operator, symmetry implies self-adjointness.
(3) Boundary Conditions: By definition, $E_{0}=\mathrm{id}$ implies $e_{\mathcal{N}(0)}=\mathbf{1}$, and $E_{1}=E$ implies $e_{\mathcal{N}(1)}=e_{\mathcal{N}}$. ∎
While the operators $e_{\mathcal{N}(s)}$ established above possess desirable analytic properties, specifically boundedness and self-adjointness, the construction suffers from a fundamental algebraic defect. For intermediate values $s\in(0,1)$, the interpolated maps fail to satisfy the idempotency condition:
$E_{s}^{2}\neq E_{s}.$ (11)
Consequently, the image $E_{s}(\mathcal{M})$ does not constitute a von Neumann subalgebra, and thus Tomiyama’s theorem cannot be invoked. This structural failure is critical: without a subalgebra structure, the associated modular objects, such as the modular conjugation $J_{E_{s}(\mathcal{M})}$ required to generate the unitary flow, are ill-defined, rendering the path physically incomplete.